OpenStudy (anonymous):

How do you convert solve these? Someone suggested me to convert them first to complex. But I don't know what to do next. Let Z1=20∠36.87 while Z2= 20∠-52.13 (Z1*Z2)/(Z1+Z2) = P.S. The answer is 14.14∠-8.13 I don't know how though.

OpenStudy (rsadhvika):

Z1*Z2 = (20*20) ∠(-52.13 + 36.87)

OpenStudy (rsadhvika):

Z1 + Z2 = 20∠36.87 + 20∠-52.13 = 20. cos(36.87) + i 20. sin (36.87) + 20. cos(-52.13) + i 20. sin (-52.13) = ?